Compressed air is used to fire a 60 g ball vertically upward from a 0.70-m-tall tube. The air exerts an upward force of 2.4 N on the ball as long as it is in the tube. Part A How high does the ball go above the top of the tube? Express your answer to two significant figures and include the appropriate units
Answer #1

The work done by the compressed gas is given by

WD=F*s

WD=2.4 * 0.70

WD = 1.687

By Conservation Of Energy Principle, the work done by the gas will be used up to achieve height by the ball, that is, to convert this energy into potential energy

WD = PE

1.68 = mgh

1.68 = 0.06 +9.81 *

h = 2.854m